Jump to: navigation, search

User Manual Basics

Here is a link to the .pdf copy of the 74" TCS user manual, v. 10.2, updated Nov. 2014.

Call out list

Telescope software issues: standby **113 (0829034869) or 9118

Electronics Issues: standby **103 (0767895131) or 9117

Mechanical Issues: standby **104 (07678951232)


The 74" TCS is written in C using XFORMS to create the Graphical User Interface (GUI). The software runs on a PC running real-time linux.

The TCS allows the control of the telescope; the acquisition and guiding (A/G) camera; and the dome (including opening and closing of the shutters and control of the windblind, as of December 2009). As of May 2016, the TCS allows powering and slewing of the telescope.

Start-up Recipe

1. Switch ON lighting circuits at entrance to dome.

2. Switch OFF all downstairs & observing floor lights (having trouble finding the switches?)

3. Check weather on the Sutherland weather website -- only proceed if windspeed < 60 km/h; humidity <90%; T_external - T_dewpoint > 1.5.

4. Check Telescope Power Control is set to PLC at top of stairs (should be the default).

5. Turn on Telescope Power. Manual: Press and hold the green button for 3 sec. under "Power" on the control panel in the warm room. TCS: take control and press box labeled "Telescope Power".

6. Open dome shutters. (TCS/ grey panel on N. pier)

7. Open mirror covers (TCS/buttons on telescope base)

8. Switch dome rotation to Auto (blue console N. pier)

9. Lower windblind (if appropriate; TCS/grey panel N. pier)

10. Switch on slow motion drive/tracking. ("SMD" switch on control panel/TCS "Tracking" button).

11. Initialise XY-slides (TCS).

Shutdown checklist

* NOTE: in bad weather, perform step 6 first, then 7, then 1 onwards.

1. Stop instrument and autoguider exposures.

2. Initialise XY-slides (TCS)

3. Switch OFF the slow motion drive/tracking ("SMD" switch on control panel/TCS "Tracking" button).

4. Park telescope at HA = 0h, Dec = -32º in SLOW mode.

5. Switch dome rotation to Manual (blue console N. pier)

6. Close mirror covers (TCS/buttons on telescope base). If windspeed > 50km/h, rotate dome 180º from wind direction before closing.

7. Close dome shutters. (TCS/ grey panel on N. pier)

8. Rotate dome to 90º azimuth (to keep rain off telescope).

9. Lower windblind (if appropriate; TCS/grey panel N. pier).

10. Switch OFF telescope power (push and hold for 3 sec. "Power" button on control panel/TCS "Telescope Power" button)

11. Switch OFF lighting circuits at exit.

12. End of run: copy data; report faults; complete [ http://www.saao.ac.za/science/observing/feedback/ feedback form], put Sutherland Observing Record in tray in warm room; for SpUpNIC or GIRAFFE, tear out 1st and 2nd copy of each completed page in triplicate log book -- keep 1st copy and put 2nd copy in end of run tray.

Instrument Focus settings

Approximate focus values for each of the instruments are listed here. Make further adjustments on target.

  GIRAFFE: ~2090
  HIPPO: ~2120 
  SHOC"  ~1900 (~1940 with focal reducer)
  SpUpNIC:   ~2200
  STE3/4: ~2090 (~ 2120 with focal reducer)


Telescope power control

The original telescope isolator switch at the top of the stairs has been replaced by a three position switch (see figure below) that should be in the "PLC" position and left that way for normal operations, day and night. Observers should not use this switch unless authorised to do so by a technician. If the switch is not found in the "PLC" position, contact the electronics technician for guidance. Telescope power is now switched on/off at a new multifunctional control panel in the warm room, described in the following section.

Warm room control panel

In April 2013, a new panel was installed in the warm room to control power to the telescope, slow motion drive and automatic dome rotation, and to indicate the status of a number of telescope subsystems (see figure below). Its functionality is described below.

Telescope power

Large green "On" and "Off" buttons are located in the top left section of the panel, with corresponding indicators that illuminate to show the current status. To operate, push and hold in for $\sim 3$ seconds. These buttons replace the power switch previously located at the top of the stairs.

There is a safety interlock preventing telescope power from being switched off unless:

  • Dome shutter is CLOSED (both "SHUTTER" indicators are off)
  • Mirror shutter is CLOSED (both "MIRROR" indicators are off)
  • RA and Dec are set to SLOW (both "FM" indicators are off)

The power OFF indicator will flash if an attempt is made to power down the telescope while the above subsystems are not in a safe state.

Dome Information

Light switches

As you enter the 1.9m building, on the wall to your right are four switches. The two close together (labelled "Light Switch Main L.T. board" and "Light Switch D.B.5") should be switched on -- these give you mains power (switching them off kills the lights, but not the power to certain vital components such as PCs). The other two switches give you incandescent and fluorescent lights in the entrance area and should be switched off if you are going to observe/take flats/do focus tests with the spectrograph.

There is a light switch at the bottom and also at the top of the steps -- again, you need the stairwell lights off (and all observing floor lights off) if you are going to do tests on the spectrograph.

There are three other lighting systems on the observing floor which might need to be switched off:

  • Dome incandescent lights -- the switch for these is on the dome wall, to the west side of the north pier (just to the right of the ``grating room door).
  • Dome fluorescent lights just below catwalk level -- these are switched on/off at the new white switches beside the electrical panel (the metal doors on the south west wall of the dome near the stair railing)
  • Dome fluorescent lights high up on the inside of the dome itself -- these can be operated from three different places:
  1. The control room, using the DOME dropdown menu on the TCS
  2. The observing floor, using the panel on the North pier. If the switch does not work the first time, press the "Request Control" button at the top of the panel and try again.
  3. The catwalk, using the button on the orange control cabinet by the dome shutters.

Current known issues (last updated Nov 2013)

Telescope pointing problems/time server error

If you encounter telescope pointing problems, check the time and date displayed on the TCS, as the cause could be the time server. If the date/time is wrong, please call the electronics technician for assistance.